OpenStudy (compassionate):

A baseball player has hit a ball. The height, y, of the ball x seconds after it is hit is given by y= -16x^2 + 80x + 3.

OpenStudy (compassionate):

So, I need to find the time it takes the ball to reach it's make height; it's max height, and the time it takes the ball to land on the ground. y= -16x^2 + 80x + 3 I can use the vertex formula to find the time of it's max height vertex = -b/2a, -80/2(16) = -80/32 Vertex or time, then is equal to 2.5 seconds. Is the first part right?

OpenStudy (compassionate):

@Miracrown , @hartnn , @ganeshie8

OpenStudy (compassionate):

@wio

OpenStudy (compassionate):

@dan815

OpenStudy (compassionate):

y= -16x^2 + 80x + 3 Take 5/2 y = -16(5/2)^2 + 80(5/2) + 3 y = -16(25/4) + 80(5/2) + 3 y = -100 + 200 + 3 y = 103 So my points are (2.5, 103)... 2.5 being my time and 103 being my max height.

Nnesha (nnesha):

i guess there is a typo in first part. you forgot the negative sign `a=-16` -80/32 would be negative 5/2 it's \[\frac{ -80 }{ 2(-16) }=\frac{ -80 }{ -32 }=\frac{ 5 }{ 2 }\]

Nnesha (nnesha):

2nd part is correct y =103 ft (2.5 , 103) is the vertex point(max) \(\huge\color{green}{\checkmark}\)

Nnesha (nnesha):

3rd part `the time it takes the ball to land on the ground` (x-intercept) pretty sure you can solve this one too. :=) goodluck!\o^_^o/